# The Geometry behind the Two-Dimensional Hard-Disk ‘Glass Transition’

### v1

@@ Appendix C Pressure
Figure 10: Determining the phase transition by measuring the pressure. (a-c) Pressure curves (blue points) showing the reduced pressure $\beta p\sigma^{2}$$\beta$ is the inverse thermal energy and $\sigma$ is the diameter of the largest particle — as a function of the inverse particle density $v=V/N=\rho^{-1}$. The relevant size ratio $R$ is provided in the bottom-left corner of the graphs. The red curves represent polynomial fits to the data. (a) This graph shows a clear van-der-Waals loop for which the coexistence pressure is indicated by the horizontal dashed black line. The two extracted coexistence points are indicated in green and the associated $v$ using vertical dotted purple lines. (b) This graph shows an inflection point, but no longer a van-der-Waals loop. (c) The $R^{-1}=1.2$ data still has a change in slope, for which the minimum slope can be fitted. (d) The $R$-$\phi$ state diagram of Fig. 4 with the $\phi$ for which there is dynamic arrest indicated in blue. On top of this, I have added the pressure data extracted from the van-der-Waals loops and changes in slope in purple. The solid purple curves guide the eye for the coexistence area fractions, while the dashed purple curve guides the eye for the minimum-slope data below $R\approx 0.87$. Below $R\approx 0.82$ I was unable to accurately determine any such minimum. Error bars denote the standard error of the mean in all cases. The cyan data points labelled “$r$” are obtained from Ref. [89], wherein coexistence $\phi$ and $p$ were determined.
@@
For sufficiently low values of $R$i.e., values of the size ratio below $R\approx 0.87$, I was unable to detect a loop structure. However, my data was still of sufficient quality to establish a minimum slope. That is, as the loop vanishes, there is an inflection point, suggestive of a critical point, see Fig. 10b. Lowering $R$ further (up to $R\approx 0.82$), I still observe a change in slope, see Fig. 10c, for which I can determine the value of $v$ (or equivalently $\phi$) where the slope attains a local minimum. Figure 10d shows the coexistence points and local minima obtained using the pressure curves on top of the state diagram from Fig. 4. I observe that for $R$ close enough to $1$ the top coexistence curve matches the pseudo-polygon curve $n_{\bar{q}}=5.5$ quite well.
@@
I also compared my coexistence data to that reported in the literature [89], which was obtained using a standard Monte Carlo simulation technique based on the Metropolis algorithm. The number of particles considered in this study is smaller (400), yet I observe an acceptable agreement between their results and mine for $R\approx 1$. For lower $R$, I do not pick upon coexistence, presumably because there is a difference in the level of equilibration. Note that obtaining coexistence was not the aim of this study nor is it easy to compare Monte Carlo and EDMD directly.
@@ Appendix D Shell and Shape Percolation
I desired to confirm the nearly constant value of $\phi_{a}$ in the range $R\lesssim 0.8$, also see Fig. 4. In order to independently measure this, I considered percolation in the system. It should be clear that when working with hard particles, contact percolation only occurs at rcp. However, when there is a substantial structural change or the system falls out of equilibrium there should be a change in the number of collisions. As the temperature — the average particle velocity — is maintained constant, physical intuition suggests that the gaps between particles should reflect such a change in collision rate. In post-processing the data, I analyzed this as follows.
@@
I augmented the size of each particle with a small shell of size $\epsilon$, that is, when particle $i$ has a diameter $\sigma_{i}$, I treat it as having a diameter $\sigma_{i}+2\epsilon$, see Fig. 11a. I use a cluster algorithm to identify the largest cluster and establish whether the system is percolating by checking whether it spans the simulation box in both directions [87, 126], see the ESI [59] for the code. This is done over all realizations of the system at a given $R$ and $\phi$ value.
@@
Figure 11: Percolation transitions based on contact and isoperimetric quotient. (a) Illustration of the shell-based percolation procedure. Each hard-disk particle (red) with diameter $\sigma$ is surrounded by a shell (blue) of width $\epsilon$, which does not impact the interactions. These shells are used to determine local clusters, as illustrated for a small cluster of four particles using green lines. Note here I only show a monodisperse example, but the procedure is equivalent for a bidisperse system. (b,c) The value of the area fraction $\phi$ for which the percolation transition occurs as a function of the shell width in log-linear representation. The dashed red line indicates the asymptotic value for $\phi$ when $\epsilon\downarrow 0$, which corresponds to the rcp value $\phi_{m}$ shown in Fig. 4. In panel (b), the system is monodisperse ($R=1$) and there appear to be two branches to the curve, as indicated using the dotted black lines, which guide the eye. In panel (c), the system is bidisperse with an inverse size ratio $R^{-1}=1.7$. Note that the curve appears relatively smooth. (d) Approach to finding the value of $\epsilon$ for which the system changes. The curve shows the numerically computed value of $(\partial\log\left[\phi_{m}-\phi\right])/(\partial\log\epsilon)$, which shows a clear peak near $\log\epsilon=-3.5$. This peak is fitted using a fourth-order local polynomial fit (red curve) and used to establish the $\epsilon$ value for which the system undergoes a change, $\epsilon^{\ast}$, as indicated using the dotted black line. (e) Representation of the $R$-$\phi$ diagram in Fig. 4, which shows the $\phi^{\ast}$ following from the shell percolation $\epsilon^{\ast}$ in magenta. The blue data shows the kinetic arrest obtained from the MSDs. Another form of percolation was considered by establishing when a system-spanning cluster of local neighborhoods with $n_{q}\geq 5$ was obtained (orange data). That is, a cluster of particles whose neighborhood is closer to being hexagonal than it is to being pentagonal. The dots indicate the data, the error bars show the standard error of the mean, and the curves guide the eye.
@@
I then use this data to establish a fraction of times that the system is clustered. Between $\phi=0$ and the rcp value $\phi_{m}$, I always find some fraction of the systems to clusters and this $\phi$ dependence can be fitted accurately using a sigmoid, see the ESI [59]. The mid-point of the sigmoid is used to establish the percolation area fraction $\phi(\epsilon)$, as shown in Fig. 11b,c. The idea is that in the limit of an infinite system size, the sigmoid becomes a Heaviside function about the percolation $\phi$.
It is clear that the percolation area fraction tends asymptotically toward the rcp value $\phi_{m}$ for $\epsilon\downarrow 0$. Additionally, as $\epsilon\uparrow\infty$, I have that $\phi(\epsilon)\downarrow 0$. Interestingly, I see that there are two branches to this shell-percolation $\phi$ when the system undergoes a (first-order) phase transition from a disordered fluid to a hexagonal crystal, see Fig. 11b for $R=1$. While the two branches are visually obvious, it proved difficult to fit the end points of the two branches with sufficient accuracy. That is, determining the associated value of $\epsilon$ and then inferring coexistence $\phi$ should be possible. However, given the number of numerical operations carried out to arrive at $\phi(\epsilon)$, it is difficult to obtain a precise $\phi$ value from an $\epsilon$ with significant variance.
The data appears more smooth when the system does not seem to undergo a first-order phase transition, see Fig. 11c for $R^{-1}=1.7$. However, given the quality of the data it is not possible to state definitively whether the curve is smooth or that there remains a small gap, as was the case for $R=1$. Nonetheless, it is possible to extract the $\phi_{a}$ from this seemingly smooth percolation line. I measured the distance to $\phi_{m}$ and take the natural logarithm of this separation $(\phi_{m}-\phi)$ as well as that of $\epsilon$.
From this data I can numerically determine the local power-law exponent $(\partial\log\left[\phi_{m}-\phi\right])/(\partial\log\epsilon)$, using a simple central differences scheme. This local power-law exponent exhibits a clear peak, see Fig. 11d, which can be fitted using a fourth-order polynomial $a+b\log^{2}\epsilon+c\log^{4}\epsilon$ (red curve). The peak position $\epsilon^{\ast}$ in turn can be used to establish the position of the peak $\phi(\epsilon^{\ast})$. Note that this position corresponds to a change in slope in the original $\phi(\epsilon)$ data, which is what I sought to locate.
@@
The corresponding area fraction is shown in Fig. 11e for those points where the was no clear indication that there were two branches (magenta data). It is directly apparent that the data accurately fits $\phi_{a}$ (blue data) for $R\lesssim 0.80$. For the remaining range, $R\gtrsim 0.81$, the trend is more close to the one of the $n_{\bar{q}}=5.5$ curve (red).
@@
Lastly, I studied whether there was any significance to shape percolation. That is, I note that crystallization is accurately predicted using $n_{\bar{q}}=5.5$, but that this is a poor predictor of $\phi_{a}$. Therefore, I sought to establish those $\phi$ for which percolation is obtained when particles are added to a cluster whenever $n_{q}\geq 5.5$. This data is shown using orange in Fig. 11e. The last bit of the guide to the eye is dashed because it was not clear whether there was a branched or a continuous $\phi(\epsilon)$ curve. For $R\gtrsim 0.83$, the curve lies half-way between the two coexistence values of $\phi$ as established using the pressure calculation in Fig. 10. These ares not shown here to avoid cluttering the figure. I conclude that there is limited value to this measure.
@@ Appendix E Median Isoperimetric Quotient
Appendix F Perspective on 3D Hard-Sphere Systems
@@
Let us attempt to apply the notion of a geometric ground state to 3D hard-sphere systems, for which the volume fraction $\eta$ is the control parameter. No experiments similar to the ones performed by Lozano et al. [1] and Li et al. [2] have been performed in 3D to the best of my knowledge. Given that Lozano et al. placed the glass transition at $\phi\approx\phi_{g}\approx 0.777$, it makes sense to try to connect to observations on the 3D glass transition. Acknowledging that the dynamics in 3D may be very different from 2D, this at least provides a reference point.
The literature places the hard-sphere glass-transition across a wide range of $\eta$ close to $\eta=0.6$. For example, Zaccarelli et al. [127] provide $\eta_{g}\approx 0.588$ for the smallest spheres in a polydisperse sample. A recent work by Geiger et al. [128] on a bidisperse sphere mixture places the transition at $\eta_{g}\approx 0.61$. Given that my 2D results are obtained for bidisperse systems, the value of $\eta_{g}=0.61$ will serve as my reference point.
I start by assuming that I can focus on monodisperse spheres that are arranged in such a way that their Voronoi cells tile space, drawing the analogy to my findings in 2D. Elements of a subclass of space-filling polyhedra, referred to as plesiohedra [129], meet this requirement. Among these, cubic Voronoi cells are obtained for a simple-cubic lattice with $\eta=\pi/6$. Truncated-octahedron-shaped Voronoi cells correspond to a body-centered cubic arrangement (bcc) with $\eta=\sqrt{3}\pi/8$. The face-centered cubic (fcc) arrangement has Voronoi cells shaped like a rhombic dodecahedron and $\eta=\pi/\sqrt{12}\approx 0.9069$. Spheres arranged in a hexagonal close packing (hcp) also have $\eta=\pi/\sqrt{12}$, but their Voronoi cell is a trapezo-rhombic dodecahedron has two types of faces.
@@
Removing the central particle out of a neighborhood in an fcc arrangement, gives a volume fraction $\eta=\pi\sqrt{12}/13\approx 0.8371$ and associated Voronoi cells that are rhombic dodecahedron with one face extended into a point. Removing a central particle in an hcp arrangement will give two types of Voronoi cell, as the base is a trapezo-rhombic dodecahedron, which has both trapezoids and rhombi as its sides. None of these elements have a volume fraction that convincingly matches $\eta_{g}$. Turning to the bcc lattice, I also gain two types of Voronoi cell by removing the ‘body’ from the body-centered cubic arrangement. There are 6 next-nearest neighbors for which the square face is extended, leading to a Voronoi cell with $\eta=6\sqrt{3}\pi/49\approx 0.666294$. Additionally, there are 4 nearest neighbors for which the hexagonal face is extended, and I find $\eta=4\sqrt{3}\pi/35\approx 0.621874$. This is relatively close to the reported value of $\eta_{g}$, but the connection is not strong. Potentially one could take the mean of the two and obtain a value $\eta\approx 0.644$ that is close to the random-close packing value for hard spheres in 3D [122]. Though this is highly speculative.
A simple cubic arrangement of particles gives $\eta=\pi/6\approx 0.52360$, which significantly underestimates the transition value, placing it closer to the prediction ($\eta\approx 0.52$) of mode-coupling theory [130]. Lastly, I consider the triakis truncated tetrahedron, which is generated by the diamond lattice. This polyhedron tiles space and gives rise to $\eta=\sqrt{3}\pi/16\approx 0.34009$. It may be a good candidate for a cage-formation ground state, drawing the analogy to my 2D observation for the honeycomb lattice, see Section VIII.4. If my geometric-ground-state framework turns out to have merit in 3D and the diamond lattice can indeed influence the fluid, this effect should already be present for unexpectedly low volume fractions. It would therefore be interesting to see what experiments similar to the ones in Ref. [2] would reveal in a 3D system.
@@
There is one final possibility for generating uniform Voronoi cells that I could find, namely to use a tetrahedral-octahedral honeycomb. In this arrangement, the particles are the inscribed spheres to the octahedra comprising the honeycomb and each of the two tetrahedra to an octahedron contributes 1/4 of its volume to the Voronoi cell. This gives the volume fraction $\eta=8\pi/(27\sqrt{3})\approx 0.537422$, which is clearly not close to the glass transition in 3D. However, as indicated in the main text, it is close to where the random-loose packing value for a frictional hard-sphere system is placed in the literature.
Concluding, I appreciate that 3D systems are significantly more complex in terms of their geometry than 2D ones. In my brief analysis of plesiohedra and tilings, I was unable to identify a particularly good candidate geometric ground state that could match $\eta_{g}$. This does not mean that there are no effects in 3D and it does not mean to imply that the glass transition in 3D needs to be understood in this framework. Further work is clearly required in this direction before any definitive conclusions can be drawn.

### v2

@@ Appendix C Crystallinity and Pressure
Figure 12: Determining the phase transition via the Steinhardt order parameter and by measuring the pressure. (a) The order parameter $\psi_{6}$ indicated in the state diagram of Fig. 6, the use of color is indicated by the legend. Only the RCP (black), the $n_{\bar{q}}=5.5$ (red), and the $\max_{\phi}\!\left(n_{\bar{q}}\right)$ (green) curves are shown from the data presented in Fig. 6. (b-c) Pressure curves (blue points) showing the reduced pressure $\beta p\sigma^{2}$$\beta$ is the inverse thermal energy and $\sigma$ is the diameter of the largest particle — as a function of the inverse particle density $v=V/N=\rho^{-1}$. The relevant size ratio $R$ is provided in the bottom-left corner of the graphs. The red curves represent polynomial fits to the data. (a) This graph shows a clear van-der-Waals loop for which the coexistence pressure is indicated by the horizontal dashed black line. The two extracted coexistence points are indicated in green and the associated $v$ using vertical dotted purple lines. (b) This graph shows an inflection point, but no longer a van-der-Waals loop. (c) The $R^{-1}=1.2$ data still has a change in slope, for which the minimum slope can be fitted. (d) The $R$-$\phi$ state diagram of Fig. 6 with the $\phi$ for which there is dynamic arrest indicated in blue. On top of this, I have added the pressure data extracted from the van-der-Waals loops and changes in slope in purple. The solid purple curves guide the eye for the coexistence area fractions, while the dashed purple curve guides the eye for the minimum-slope data below $R\approx 0.87$. Below $R\approx 0.82$ I was unable to accurately determine any such minimum. Error bars denote the standard error of the mean in all cases. The cyan data points labelled “s” are obtained from Ref. [120], wherein coexistence $\phi$ and $p$ were determined.
In the canonical ensemble, there is a van-der-Waals (vdW) loop in the pressure $p$ as a function of the inverse particle density $v=V/N=\rho^{-1}$. I therefore computed the pressure from the collisions in the system via the momentum exchange during collisions, using the standard virial expression. This was averaged and subsequently used to establish vdW loops whenever they are present, see Fig. 12b for an example of this structure. For sufficiently low values of RRi.e., values of the size ratio below $R\approx 0.87$, I was unable to detect a loop structure. However, my data was still of sufficient quality to establish a minimum slope. That is, as the loop vanishes, there is an inflection point, suggestive of a critical point, see Fig. 12b.
@@
Lowering $R$ further (up to $R\approx 0.82$), I still observe a change in slope, see Fig. 12c, for which I can determine the value of $v$ (or equivalently $\phi$) where the slope attains a local minimum. Figure 12d shows the coexistence points and local minima obtained using the pressure curves on top of the state diagram from Fig. 6. I observe that for $R$ close enough to $1$ the top coexistence curve matches the pseudo-polygon curve $n_{\bar{q}}=5.5$ quite well. The agreement is less convincing for lower $R$, but this is an additional piece of information rather than the main goal of the study. Given the intrinsic difference in ensemble, no attempt was made to improve the quality of the data.
@@
Lastly, I compared my coexistence data to that reported in the literature [120], which was obtained using a standard Monte Carlo simulation technique based on the Metropolis algorithm. The number of particles considered in this study is smaller (400), yet I observe an acceptable agreement between their results and mine for $R\approx 1$. For lower $R$, I do not pick upon coexistence, presumably because there is a difference in the way the studies are set up.
@@ Appendix D Shell and Shape Percolation
In order to confirm the nearly constant value of $\phi_{a}$ in the range $R\lesssim 0.8$, also see Fig. 6, I considered percolation in the system. When working with hard particles, contact percolation only occurs at RCP. However, when there is a substantial structural change, or the system falls out of equilibrium, there should be a trace of this in the number of collisions. As the temperature — the average particle velocity — is maintained constant, physical intuition suggests that the gaps between particles should reflect such a change in collision rate. This was analyzed as follows.
@@
The size of each particle was augmented with a small shell of size $\epsilon$, that is, when particle $i$ has a diameter $\sigma_{i}$, it is treated as having a diameter $\sigma_{i}+2\epsilon$, see Fig. 13a. Next, I used a cluster algorithm to identify the largest cluster and establish whether the system is percolating by checking whether it spans the simulation box in both directions [118, 164], see the ESI [64] for the code. This is done over all realizations of the system at a given $R$ and $\phi$.
@@
Figure 13: Percolation transitions based on contact and isoperimetric quotient. (a) Illustration of the shell-based percolation procedure. Each hard-disk particle (red) with diameter $\sigma$ is surrounded by a shell (blue) of width $\epsilon$, which does not impact the structure. These virtual shells are used to determine local clusters, as illustrated for a small cluster of four particles using green lines. Only a monodisperse example is shown, but the procedure is equivalent for a bidisperse system. (b,c) The value of the area fraction $\phi$ for which the percolation transition occurs, as a function of the shell width $\epsilon$ in log-linear representation. The dashed red line indicates the asymptotic value for $\phi$ when $\epsilon\downarrow 0$, which corresponds to the RCP value $\phi_{m}$ shown in Fig. 6. In panel (b), the system is monodisperse ($R=1$) and there appear to be two branches to the curve, as indicated using the dotted black lines, which guide the eye to $\phi=0.699$ and $\phi=0.724$; the region over which there are transitions [121, 122]. In panel (c), the system is bidisperse with an inverse size ratio $R^{-1}=1.7$. Note that the curve appears relatively smooth. (d) Approach to finding the value of $\epsilon$ for which the system changes. The curve shows the numerically computed value of $(\partial\log\left[\phi_{m}-\phi\right])/(\partial\log\epsilon)$, which shows a clear peak near $\log\epsilon=-3.5$. This peak is fitted using a fourth-order local polynomial fit (red curve) and used to establish the $\epsilon$ value for which the system undergoes a change, $\epsilon^{\ast}$, as indicated using the dotted black line. This corresponds to an area fraction of $\phi\approx 0.777$. (e) Representation of the $R$-$\phi$ diagram in Fig. 6, which shows the $\phi^{\ast}$ following from the shell percolation $\epsilon^{\ast}$ in magenta. The blue data shows the kinetic arrest obtained from the MSDs. Another form of percolation was obtained by establishing when a system-spanning cluster of local neighborhoods with $n_{q}\geq 5$ was obtained (orange data and solid orange curve). That is, a cluster of particles whose neighborhood is closer to being hexagonal than it is to being pentagonal. The dots indicate the data, the error bars show the standard error of the mean, and the curves guide the eye.
@@
Using this data, the fraction of times that the system is clustered can be established. Between $\phi=0$ and the RCP value $\phi_{m}$, I always found some fraction of the systems to clusters and this $\phi$-dependence can be fitted accurately using a sigmoid  [^14] In the limit of an infinite system size, the sigmoid becomes a Heaviside function about the percolation $\phi$., see the ESI [64]. The mid-point of the sigmoid is used to establish the percolation area fraction $\phi(\epsilon)$, as shown in Fig. 13b,c; these are two representative samples of the full data.
It is clear that the percolation area fraction tends asymptotically toward the RCP value $\phi_{m}$ for $\epsilon\downarrow 0$. Interestingly, there are two branches to this shell-percolation $\phi$, when the system undergoes a (first-order) phase transition from a disordered fluid to a hexagonal crystal, see Fig. 13b for $R=1$. While the two branches are visually obvious, it proved difficult to fit the end points of the two branches with sufficient accuracy. The correspondence with known phase transition results is therefore only indicated visually in Fig. 13b. The data appears smooth when the system does not seem to undergo a first-order phase transition, see Fig. 13c for $R^{-1}=1.7$. For such data sets, it proved possible to extract the $\phi_{a}$ by measuring the distance to $\phi_{m}$ and take the natural logarithm ($\log$) of this separation $(\phi_{m}-\phi)$ as well as that of $\epsilon$. From this data I could numerically determine the local power-law exponent $(\partial\log\left[\phi_{m}-\phi\right])/(\partial\log\epsilon)$ using central differences. This local power-law exponent exhibits a clear peak, see Fig. 13d, which was fitted using a fourth-order polynomial $a+b\left(\log\epsilon\right)^{2}+c\left(\log\epsilon\right)^{4}$ (red curve).
@@
The peak position $\epsilon^{\ast}$ in turn was used to establish the equivalent area fraction $\phi(\epsilon^{\ast})$. Note that this position corresponds to a change in slope in the original $\phi(\epsilon)$ data. The corresponding area fraction is shown in Fig. 13e for those points where the was no clear indication that there were two branches (magenta data). It is directly apparent that the data accurately fits $\phi_{a}$ (blue data) for $R\lesssim 0.80$. For the remaining range, $R\gtrsim 0.81$, the trend is closer to the one of the $n_{\bar{q}}=5.5$ curve (red).
@@
Lastly, I studied whether there was any significance to shape percolation. That is, I note that crystallization is accurately predicted using $n_{\bar{q}}=5.5$, but that this is a poor predictor of $\phi_{a}$. Therefore, I sought to establish those $\phi$ for which percolation is obtained when particles are added to a cluster whenever $n_{q}\geq 5.5$. This data is shown using orange in Fig. 13e. The last bit of the guide to the eye is dashed because it was not clear whether there was a branched or a continuous $\phi(\epsilon)$ curve. For $R\gtrsim 0.83$, the curve lies half-way between the two coexistence values of $\phi$ as established using the pressure calculation in Fig. 12. These are not shown here to avoid cluttering the figure. I conclude that there is limited value to this measure.
@@ Appendix E Median Isoperimetric Quotient
Appendix F Fluctuations of the Particle Neighborhood
In this appendix, the frequency by which particle neighborhoods changes into one another is analyzed. The goal of this is to connect to the observations of Narinder et al. [66], where structure and fluctuations of a length-wise polydisperse fluid of colloidal rods was reported. In brief, for each time step, I computed the radical Voronoi diagram and kept track of the $q$ value belonging to each particle. I found that for sufficiently high $\phi$, the neighborhoods dwell about integer values of $n_{q}$ and the system does not favor half-integer $n_{q}$. For example, the trajectory of a certain particle may lead it to explore pentagonal ($n_{q}\approx 5$) and hexagonal ($n_{q}\approx 6$) neighborhoods, but the particle spends little time in $5.5$-gonal neighborhoods  [^15] This observation does not conflict with the fact that in Fig. 4a-c, I observe dips in the $P_{\phi}(q)$ about integer values of $n_{q}$ and a smooth increasing trend near half-integer values of $n_{q}$. There is a bias to the representation of the system using $P_{\phi}(q)$, as explained in Section IV.. Next, I extracted transition rates for the exchange between neighborhoods and normalized this by the total number of particle displacements in the system. This resulted in, for example, relative exchange rates between pentagonal and hexagonal neighborhoods, $r_{5\leftrightarrow 6}$, some which are shown in Fig. 15.
Figure 15: The relative neighborhood exchange rate indicates strong fluctuations close to transitions. The ratio of the number of changes between a pentagonal and a hexagonal neighborhood to the total number of particle displacements in the system, $r_{5\leftrightarrow 6}$, as a function of the area fraction $\phi$. The red and blue points show data gathered for $R=1$ and $R^{-1}=1.4$, respectively, as labelled. The solid, light-red and light-blue curves serve to guide the eye. The peak positions are indicated by the magenta and purple vertical dashed lines, respectively. Errors are smaller than the symbol size and therefore not shown.
Surprisingly, the relative number of exchanges constitutes a substantial fraction of particle displacements even at high $\phi$ close to $\phi_{a}$ or equivalently $\phi_{g}$. This means that while there is relatively limited cage hopping, as evidenced by the low value of $D(\phi)$, the cages themselves are quite dynamic. Additionally, I find no significant difference in the $r_{i\leftrightarrow i+1}$ between small and large particles for any integer $i$, see the ESI [64]. This implies that the cage dynamics is shared equally among the large and small particles, though it does not exclude a difference in cage-hopping rate, i.e., in the particles’ respective diffusivity.
Both data sets presented in Fig. 15 show a peak. When there is crystallization, as is the case for $R=1$, it is sharp and very pronounced. Fitting the peak values results in $\phi\approx 0.703$ and $\phi\approx 0.757$ for $R=1$ and $R^{-1}=1.4$, respectively. The former agrees very well with where the fluid state is expected to become unstable. The latter is close to, but not quite at $\phi_{g}$. For $R^{-1}=1.4$, the peak is also broader and less pronounced. Within the error, I detected no discernable drift toward one type of neighborhood as a function of time.
That the rate of cage rearrangement is relatively high, fits with the conceptual picture of a self-propelled probe particle (experiments in Ref. [1, 66]) picking up on the fluctuations of the local environment, as put forward in recent modeling work from my group [65, 66]. The fact that $r_{5\leftrightarrow 6}$ is asymmetric about $\phi\approx\phi_{g}$ would also fit with the asymmetric probe response observed in Ref. [1]. However, there is a slight mismatch between the peak value of $\phi$ and $\phi_{g}$. This should be addressed in future work, wherein self-propelled probes are included in the sample, but this falls outside of the scope of the present research.
Appendix G Perspective on 3D Hard-Sphere Systems
@@
Lastly, I will attempt to apply the notion of a geometric ground state to 3D hard-sphere systems, for which the volume fraction $\eta$ is the control parameter. To the best of my knowledge, no experiments similar to the ones performed by Lozano et al. [1] and Li et al. [4] have been performed in 3D. Given that Lozano et al. placed the glass transition at $\phi\approx\phi_{g}\approx 0.776$, it makes sense to try to connect to observations on the 3D glass transition, as well as 3D jamming in view of the 2D experiments of Refs. [82, 2, 3]. Acknowledging that the dynamics in 3D may be very different from 2D, this at least provides a reference point.
For jamming of 3D granular matter, the transition value is dependent on the details of the interaction. For frictionless, monodisperse hard spheres, the value for the RCP (frictionless jamming) is typically reported to be $\eta_{j}\approx 0.64$ [167, 168, 169], while for frictional spheres, the jamming point lies at $\eta_{j}\approx 0.555$ [144, 135, 75, 145]. Singh et al. [81] studied 3D jamming for no ($\eta_{j}\approx 0.648$), sliding ($\eta_{j}\approx 0.570$), and sliding & rolling ($\eta_{j}\approx 0.365$) friction using numerical simulations and obtained three branches. It is clear from these numbers that making the connection between dynamical arrest in a Brownian suspension and RLP in a granular system is considerably more tentative.
Nonetheless, assuming that it is possible to focus on monodisperse spheres, the following analogy to the 2D hard-disk findings reported in the main text can be drawn. These should be arranged in such a way that their Voronoi cells tile space, are congruent, and result from removing particles from the closest crystalline packing via T2 transitions. The closest packing is the face-centered cubic (FCC) arrangement, which has Voronoi cells shaped like a rhombic dodecahedron and $\eta_{c}=\pi/\sqrt{18}\approx 0.740480$. However, spheres may also be arranged in a hexagonal close packing (HCP), which has an identical packing fraction $\eta=\eta_{c}$, but the associated Voronoi cell is a trapezo-rhombic dodecahedron has two types of faces. It is well known that the FCC lattice is thermodynamically preferred over the HCP lattice, although the difference in free energy per particle between the two is small [170].
@@
Removing the central particle out of a neighborhood in an FCC arrangement, in the same spirit as done for the creation of the FPT in 2D, gives a volume fraction $\eta=12\eta_{c}/13=2\sqrt{2}\pi/13\approx 0.683520$ and associated Voronoi cells that are rhombic dodecahedron with one face extended into a point. Removing a central particle in an HCP arrangement will give two types of Voronoi cell, as the base element for HCP is a trapezo-rhombic dodecahedron, which has both trapezoids and rhombi as its sides. Both of these cells have a local volume fraction identical to that of the FCC arrangement and do not provide additional possibilities for matching. Potentially, this lack of uniform congruency upon T2 removal speaks in favor of the stability of the FCC arrangement over the HCP arrangement, but this is speculative. Turning to the body-centered cubic (BCC) lattice, I also gain two types of Voronoi cell by removing the ‘body’ from the body-centered cubic arrangement. There are 6 next-nearest neighbors for which the square face is extended, leading to a Voronoi cell with $\eta=6\sqrt{3}\pi/49\approx 0.666294$. Additionally, there are 4 nearest neighbors for which the hexagonal face is extended, and I find $\eta=4\sqrt{3}\pi/35\approx 0.621874$. Taking the mean of the two leads to a value $\eta\approx 0.644$, which is close to the random-close packing value for hard spheres in 3D [146]. Though this seems hard to justify, also given that generally RCP values seem to strongly depend on $\phi$ (in 2D) and $\eta$ (in 3D).
Relaxing the requirement for removing particles, gives another congruent 3D tiling. This is generated by the diamond lattice and its Voronoi cells are triakis truncated tetrahedra. This lattice is associated with a volume fraction $\eta=\sqrt{3}\pi/16\approx 0.34009$. It may be a good candidate for a form of jamming point, drawing the analogy to my 2D observation for the honeycomb lattice, see Section VIII.2. In view of the work of Singh et al. [81], it may be a candidate geometric ground state for particles with both sliding and rolling friction. Although, the proposed value based on this argument underestimates the jamming point. It may be that this is because the analogy is not applicable, or that there is slippage in the simulations. The connection, if any, should be investigated more closely.
@@
There are several other possibilities for generating uniform Voronoi cells. Among these, the use of a tetrahedral-octahedral honeycomb has the greatest potential. In this arrangement, the particles are the inscribed spheres to the octahedra comprising the honeycomb and each of the two tetrahedra to an octahedron contributes 1/4 of its volume to the Voronoi cell. This gives the volume fraction $\eta=8\pi/(27\sqrt{3})\approx 0.537422$, which is clearly not close to the glass transition in 3D. However, as also indicated in the main text, it is close to where the random-loose packing value for a frictional hard-sphere system is placed in the literature. In contrast, the honeycomb that results from the Bilinski dodecahedron has an approximate volume fraction of $\eta\approx 0.247$, the one that follows from the ten-of-diamonds decahedron has $\eta\approx 0.393$, and the one from the acute golden rhombohedron has $\eta\approx 0.424$. None is close to a known instance of dynamical arrest or jamming, and other polyhedral space fillers give similarly low volume fractions. Certainly, with the exception of the RLP value, the proposed relation is less convincing than it is in 2D.